OpenStudy (anonymous):

How do I solve n/5-3n/10=1/5? Please Help!!!!!

OpenStudy (anonymous):

Do you mean this? \[ \frac{ n }{ 5 } - \frac{ 3n }{ 10 } = \frac{ 1 }{ 5 }\]

OpenStudy (anonymous):

Yes!!!!!

OpenStudy (anonymous):

Alright then, let's break it down.

OpenStudy (anonymous):

What we should do here, is multiply the equation out by 10.

OpenStudy (anonymous):

That way, we get the denominators out of the way.

OpenStudy (anonymous):

So we should get this:

OpenStudy (anonymous):

\[2n - 3n = 2\]

OpenStudy (anonymous):

Now its just simple algebra.

OpenStudy (anonymous):

Thank you soooo much!!!!

OpenStudy (anonymous):

No problem, happy to help.